# Weak coloring numbers of minor-closed graph classes

### v1

@@ 7.2. The base case
Let $C$ be a component of $G-S_{0}$. By 42.(b) and Lemma 37, there exists a family $\mathcal{B}_{C}$ of at most $2\left(d(h-1)+\binom{h-1}{2}\right)-1\leqslant 2dh^{2}$ bags of $\mathcal{D}$ such that $N_{G}(V(C))\subseteq\bigcup\mathcal{B}_{C}$ and for every component $C^{\prime}$ of $G-\bigcup\mathcal{B}_{C}$, $N_{G}(V(C^{\prime}))$ intersects at most two components of $G-V(C^{\prime})$. Let $S_{C}=V(C)\cap\bigcup\mathcal{B}_{C}$. Then, for every component $C^{\prime}$ of $G-(S_{0}\cup S_{C})$ intersecting $V(C)$, $C$ is a component of $G-\bigcup\mathcal{B}_{C}$, and so $N_{G}(V(C^{\prime}))$ intersects at most two component of $G-V(C^{\prime})$.
@@
$\leqslant\bar{c}_{0}(h,d)k\cdot\log r+2dh^{2}k$ by 42.(c)
$\leqslant\left(\bar{c}_{0}(h,d)+2dh^{2}\right)k\cdot\log r.$
@@ 7.3. Induction
Let $t$ be positive integers with $t\geqslant 2$. Let $X$ be a graph with $\operatorname{rtd}_{2}(X)\leqslant t$. There exist an integer $\bar{c}(t,X)$ such that for every integer $k$, for every graph $G$ with $\operatorname{tw}(G)<k$, for every family $\mathcal{F}$ of nonnull connected subgraphs of $G$, if $G$ has no $\mathcal{F}$-rich model of $X$, then there exists $S\subseteq V(G)$ such that
@@
We proceed by induction on $t$. When $t=2$, by Observation 21, $X$ is a forest. Let $h,d$ be positive integers such that $X\subseteq F_{h,d}$, and let $\bar{c}_{0}(h,d)$ be the constant given by Lemma 44. The assertion with $\bar{c}(t,X)=\bar{c}_{0}(h,d)+2dh^{2}$ follows by applying Lemma 44. Next, let $t\geqslant 3$, and assume that the result holds for $t-1$. We refer to this property as the main induction hypothesis.
@@
Let $Y$ be a graph with $\operatorname{rtd}_{2}(Y)\leqslant t-1$. There is an integer $\bar{c}_{1}(t,Y)$ such that for every positive integer $k$, for every connected graph $G$ with $\operatorname{tw}(G)<k$, for every nonempty set $U$ of vertices of $G$ with $|U|\leqslant 2$, for every family $\mathcal{F}$ of nonnull connected subgraphs of $G$, if $G$ has no $\mathcal{F}$-rich model of $K_{1}\oplus Y$, then there exist $S\subseteq V(G)$, a tree $T$ rooted in $s\in V(T)$, and a tree partition $\big{(}T,(P_{x}\mid x\in V(T))\big{)}$ of $G[S]$ with $P_{s}=U$ such that
@@
    for every $x\in V(T)$ let $T_{x}$ be the subtree of $T$ rooted in $x$, then
    $\operatorname{wcol}_{r}\left(G\left[\bigcup_{y\in V(T_{x})}P_{y}\right],P_{x}\right)\leqslant\bar{c}_{1}(t,Y)k\cdot r^{t-3}\log r$
    for every integer $r$ with $r\geqslant 2$.
Proof of the claim. Let $X^{\prime}=K_{1}\sqcup Y$. By Lemma 38 applied to $k=2$ and $X^{\prime}$, there exists a graph $X^{\prime\prime}$ with $\operatorname{rtd}_{2}(X^{\prime\prime})=\operatorname{rtd}_{2}(X^{\prime})=t$ such that for every covering $S_{1},S_{2}$ of $V(X^{\prime\prime})$ there exists $i\in\{1,2\}$ such that $X^{\prime\prime}$ contains an $\{H\subseteq X^{\prime\prime}\mid\text{$H$ connected and }S_{i}\cap V(H)\neq\emptyset\}$-rich model of $X^{\prime}$. Let $\bar{c}_{1}(t,Y)=2+\bar{c}(t-1,X^{\prime\prime})$.
@@
Assume that $G-U$ is not connected. Let $\mathcal{C}$ be the family of all the components of $G-U$. Consider a component $C\in\mathcal{C}$. By induction hypothesis applied to $G_{C}=G[V(C)\cup U]$, $U$, and $\mathcal{F}|_{G_{C}}$, there exist $S_{C}\subseteq V(G_{C})$, a tree $T_{C}$ rooted in $s_{C}\in V(T_{C})$, and a tree partition $(T_{C},(P_{C,x}\mid x\in V(T_{C})))$ of $S_{C}$ in $G_{C}$ with $P_{C,s_{C}}=U$ such that
  1. (a’)
    $V(F)\cap S_{C}\neq\emptyset$ for every $F\in\mathcal{F}|_{G[V(C)\cup U]}$;
  2. (b’)
  3. (c’)
  4. (d’)
    for every $x\in V(T_{C})$ let $T_{C,x}$ be the subtree of $T_{C}$ rooted in $x$, then
    $\operatorname{wcol}_{r}\left(G_{C}\left[\bigcup_{y\in V(T_{C,x})}P_{C,y}\right],P_{C,x}\right)\leqslant\bar{c}_{1}(t,Y)k\cdot r^{t-3}\log r$
    for every integer $r$ with $r\geqslant 2$.
Then let $S=\bigcup_{C\in\mathcal{C}}S_{C}$, let $T$ be the tree obtained from the disjoint union of all the $T_{C}$ for $C\in\mathcal{C}$ by identifying all the vertices in $\{s_{C}\mid C\in\mathcal{C}\}$ into a single vertex $s$. Finally, let $P_{s}=U$ and $P_{x}=P_{C,x}$ for every $C\in\mathcal{C}$ and $x\in V(T_{C})-\{s_{C}\}$. Then we claim that (a)-(d) hold. Indeed, for every $F\in\mathcal{F}$, $F$ is connected so either $F$ is a subgraph of $G-U$ or $F$ intersects $U$. In both cases, we see that $V(F)\cap S\neq\emptyset$, so (a) holds. Item (b) holds by (b’) as $U\subseteq S_{C}$ for all $C\in\mathcal{C}$. Item (c) holds as every component of $G-S$ is a component of $G_{C}-S_{C}$ for some $C\in\mathcal{C}$. Finally, for all $x\in V(T)$ with $x\neq s$ item (d) follows directly from (d’) and the construction of $T$, while when $x=s$, $\operatorname{wcol}_{r}\left(G\left[\bigcup_{y\in V(T)}P_{y}\right],P_{s}\right)=|U|\leqslant 2$ for every integer $r$ with $r\geqslant 2$. From now on we assume that $G-U$ is connected.
Let $\mathcal{F}^{\prime}$ be the family of all the nonnull connected subgraphs $H$ of $G-U$ such that $U\cap N_{G}(V(H))\neq\emptyset$ and $F\subseteq H$ for some $F\in\mathcal{F}$. We argue that there is no $\mathcal{F}^{\prime}$-rich model of $X^{\prime\prime}$ in $G-U$. Suppose to the contrary that there is an $\mathcal{F}^{\prime}$-rich model $\big{(}B_{x}\mid x\in V(X^{\prime\prime})\big{)}$ of $X^{\prime\prime}$ in $G-U$. For each $u\in U$, let $S_{u}=\{x\in V(X^{\prime\prime})\mid u\in N_{G}(B_{x})\}$. Since the model is $\mathcal{F}^{\prime}$-rich, $\{S_{u}\}_{u\in U}$ is a covering of $V(X^{\prime\prime})$. Therefore, there exists $u\in U$ such that $X^{\prime\prime}$ contains an $\{H\subseteq X^{\prime\prime}\mid\text{$H$ connected and }S_{u}\cap V(H)\neq\emptyset\}$-rich model $\mathcal{M}$ of $X^{\prime}$. Moreover, by definition, each branch set of $\mathcal{M}$ contains a member of $\mathcal{F}$, and furthermore, $\mathcal{M}$ is a model of $X^{\prime}$ in $G$. Recall that $X^{\prime}=K_{1}\sqcup Y$. If $v$ is the vertex of $K_{1}$ is $X^{\prime}$, then the model obtained from $\mathcal{M}$ by replacing the branch set $C_{v}$ corresponding to $v$ by $C_{v}\cup\{u\}$ yields an $\mathcal{F}$-rich model of $K_{1}\oplus Y$ in $G$, which is a contradiction. This proves that there is no $\mathcal{F}^{\prime}$-rich model of $X^{\prime}$ in $G-\{u\}$.
@@
    for every $x\in V(T_{C})$ let $T_{C,x}$ be the subtree of $T_{C}$ rooted in $x$, then
    $\operatorname{wcol}_{r}\left(G_{C}\left[\bigcup_{y\in V(T_{C,x})}P_{C,y}\right],P_{C,x}\right)\leqslant\bar{c}_{1}(t,Y)k\cdot r^{t-3}\log r$
    for every integer $r$ with $r\geqslant 2$.
@@
Since for every $C\in\mathcal{C}_{1}$, $U\cap N_{G}(V(C))=\emptyset$, every edge in $G[S]$ containing a vertec in $U$ has another endpoint in $U\cup S_{0}=P_{s}\cup P_{s^{\prime}}$. Consider an edge $vw$ in $G[S]$ such that $v\in S_{0}$ and $w\in S_{C}$ for some $C\in\mathcal{C}_{1}$. Since $\big{(}T_{C},(P_{C,x}\mid x\in V(T_{C}))\big{)}$ is a tree partition of $G_{C}[S_{C}]$ with $P_{C,s_{C}}=U_{C}$ and $S_{0}\subseteq V((G-U)-V(C))$, we conclude that $w\in P_{x}$ for some $x\in V(T_{C})$ such that $s^{\prime}x$ is an edge in $T$. Finally, for every edge $vw$ of $G[S]$ with $v,w\not\in U\cup S_{0}$, $vw$ is an edge of $G[S_{C}-U_{C}]$ for some component $C\in\mathcal{C}_{1}$, and so $v\in P_{C,x}$ and $w\in P_{C,y}$ for adjacent or identical vertices $x,y$ of $T_{C}$. Then $v\in P_{x}$ and $w\in P_{y}$. It follows that $\big{(}T,(P_{x}\mid x\in V(T))\big{)}$ is a tree partition of $G[S]$.
@@
Finally, we argue (d). Let $r$ be an integer with $r\geqslant 2$ and let $x\in V(T)$. For $x=s$, $|P_{s}|=|U|\leqslant 2$, thus the assertion is clear. For $x=s^{\prime}$, we have $T_{s^{\prime}}=V(G)-U$. By (C’),
$\operatorname{wcol}_{r}(G-U,S_{0})\leqslant\bar{c}(t-1,X^{\prime\prime})k\cdot r^{t-3}\log r\leqslant\bar{c}_{1}(t,Y)k\cdot r^{t-3}\log r.$
For $x\in V(T_{C}-\{s_{C}\})$ for some $C\in\mathcal{C}_{1}$, we have $T_{x}=T_{C,x}$, thus, the asserted inequality follows from (d”). This ends the proof of the claim. $\lozenge$
@@
Let $Y$ be a graph with $\operatorname{rtd}_{2}(Y)\leqslant t-1$. There are integers $\bar{c}_{2}(t,Y)$ such that for every positive integer $k$, for every graph $G$ with $\operatorname{tw}(G)<k$, for every family $\mathcal{F}$ of nonnull connected subgraphs of $G$, if $G$ has no $\mathcal{F}$-rich model of $K_{1}\oplus Y$, then there exist $S\subseteq V(G)$ such that
@@
Proof of the claim. Let $\bar{c}_{2}(t,Y)=2\cdot\bar{c}_{1}(t,Y)$ and let $r$ be an integer with $r\geqslant 2$. We apply 4 with an arbitrary singleton $U$ in $G$ to obtain $S\subseteq V(G)$, a tree $T$ rooted in $s\in V(T)$, and a tree partition $(T,(P_{x}\mid x\in V(T)))$ of $S$ in $G$ with $P_{s}=\{u\}$ such that
@@
    for every $x\in V(T)$ let $T_{x}$ be the subtree of $T$ rooted in $x$, then
    $\operatorname{wcol}_{r}\left(G\left[\bigcup_{y\in V(T_{x})}P_{y}\right],P_{x}\right)\leqslant\bar{c}_{1}(t,Y)k\cdot r^{t-3}\log r$
    for every integer $r$ with $r\geqslant 2$.
@@
$|\operatorname{WReach}_{r}[G,S,\sigma,u]\cap P_{y}|\leqslant\operatorname{wcol}_{r}\left(G\left[\bigcup_{z\in V(T_{y})}P_{z}\right],P_{y}\right)\leqslant\bar{c}_{1}(t,X)k\cdot r^{t-2}\log r.$
@@
This concludes the proof of (c), and the claim follows. $\lozenge$
@@
Let $Y$ be a graph with $\operatorname{rtd}_{2}(Y)\leqslant t-1$ and let $h,d$ be positive integers. There exists an integer $\bar{c}_{3}(Y,h,d)$ such that for every positive integer $k$, for every graph $G$ with $\operatorname{tw}(G)<k$, for every family $\mathcal{F}$ of nonnull connected subgraphs of $G$, if $G$ has no $\mathcal{F}$-rich model of $T^{\prime}_{h,d}(Y)$, then there exists $S\subseteq V(G)$ such that
@@
Proof of the claim. We proceed by induction on $h$. When $h=1$, $T^{\prime}_{h,d}(Y)=K_{1}\oplus(Y\sqcup Y)$ and the result follows from the previous claim applied to $Y\sqcup Y$ (note that $\operatorname{rtd}_{2}(Y\sqcup Y)=\operatorname{rtd}_{2}(Y)\leqslant t-1$) by setting $\bar{c}_{3}(Y,1,d)=\bar{c}_{2}(t,Y\sqcup Y)$.
@@
$\lozenge$
Finally, by Lemma 28, for every for every graph $X$ with $\operatorname{rtd}_{2}(X)\leqslant t$, there exists a graph $Y$ with $\operatorname{rtd}_{2}(Y)\leqslant t-1$ and positive integers $h,d$ such that $X\subseteq T_{h,d}(Y)\subseteq T^{\prime}_{h,d}(Y)$. Let $G$ be a graph with $\operatorname{tw}(G)<k$ and let $\mathcal{F}$ be a family of nonnull connected subgraphs of $G$. Suppose that $G$ has no $\mathcal{F}$-rich model of $X$. By 6, there exists $S_{0}\subseteq V(G)$ such that
@@
Item (A) follows from 6.(a). Let $C^{\prime}$ be a component of $G-S$, and let $C$ be the component of $G-S_{0}$ intersecting $V(C^{\prime})$. Then $N_{G}(V(C^{\prime}))\subseteq N_{G}(V(C))\cup\bigcup_{Q\in\mathcal{Q}_{C}}V(Q)$, and since $\bigcup_{Q\in\mathcal{Q}_{C}}V(Q)$ induces a connected subgraph of $C$ and contains a neighbor of every component of $G-V(C)$ having a neighbor in $V(C)$, we deduce that $N_{G}(V(C^{\prime}))$ intersects at most one component of $G-S$. This proves that (B) holds. And finally, (C) is true once we set $\bar{c}(t,X)=\bar{c}_{3}(Y,h,d)+(4h-1)\cdot 3$ since

### v2

@@ 7.2. The base case
Let $C$ be a component of $G-S_{0}$. By 42.(b) and Lemma 37, there exists a family $\mathcal{B}_{C}$ of at most $2\left(2d(h-1)+2\binom{h-1}{2}\right)-1\leqslant 4dh^{2}$ bags of $\mathcal{D}$ such that $N_{G}(V(C))\subseteq\bigcup\mathcal{B}_{C}$ and for every component $C^{\prime}$ of $G-\bigcup\mathcal{B}_{C}$, $N_{G}(V(C^{\prime}))$ intersects at most two components of $G-V(C^{\prime})$. Let $S_{C}=V(C)\cap\bigcup\mathcal{B}_{C}$. Then, for every component $C^{\prime}$ of $G-(S_{0}\cup S_{C})$ intersecting $V(C)$, $C$ is a component of $G-\bigcup\mathcal{B}_{C}$, and so $N_{G}(V(C^{\prime}))$ intersects at most two component of $G-V(C^{\prime})$.
@@
$\leqslant\bar{c}_{0}(h,d)k\cdot\log r+4dh^{2}k$ by 42.(c)
$\leqslant\left(\bar{c}_{0}(h,d)+4dh^{2}\right)k\cdot\log r.$
@@ 7.3. Induction
Let $t$ be positive integers with $t\geqslant 2$. Let $X$ be a graph with $\operatorname{rtd}_{2}(X)\leqslant t$. There exist an integer $\bar{c}(t,X)$ such that for every integer $k$, for every graph $G$ with $\operatorname{tw}(G)<k$, for every family $\mathcal{F}$ of connected subgraphs of $G$, if $G$ has no $\mathcal{F}$-rich model of $X$, then there exists $S\subseteq V(G)$ such that
@@
We proceed by induction on $t$. When $t=2$, by Observation 21, $X$ is a forest. Let $h,d$ be positive integers such that $X\subseteq F_{h,d}$, and let $\bar{c}_{0}(h,d)$ be the constant given by Lemma 44. The assertion with $\bar{c}(t,X)=\bar{c}_{0}(h,d)+4dh^{2}$ follows by applying Lemma 44. Next, let $t\geqslant 3$, and assume that the result holds for $t-1$. We refer to this property as the main induction hypothesis.
@@
Let $Y$ be a graph with $\operatorname{rtd}_{2}(Y)\leqslant t-1$. There is an integer $\bar{c}_{1}(t,Y)$ such that for every positive integer $k$, for every connected graph $G$ with $\operatorname{tw}(G)<k$, for every nonempty set $U$ of vertices of $G$ with $|U|\leqslant 2$, for every family $\mathcal{F}$ of connected subgraphs of $G$, if $G$ has no $\mathcal{F}$-rich model of $K_{1}\oplus Y$, then there exist $S\subseteq V(G)$, a tree $T$ rooted in $s\in V(T)$, and a tree partition $\big{(}T,(P_{x}\mid x\in V(T))\big{)}$ of $G[S]$ with $P_{s}=U$ such that
@@
    for every $x\in V(T)$,
    $\operatorname{wcol}_{r}\left(G_{x},P_{x}\right)\leqslant\bar{c}_{1}(t,Y)k\cdot r^{t-3}\log r$
    for every integer $r$ with $r\geqslant 2$, where, for $T_{x}$ being the subtree of $T$ rooted in $x$, $G_{x}$ is the subgraph of $G$ induced by $U_{x}=\bigcup_{y\in V(T_{x})}P_{y}$ and the vertex sets of all the components of $G-S$ having a neighbor in $U_{x}$.
Proof of the claim.
Let $X^{\prime}=K_{1}\sqcup Y$. By Lemma 38 applied to $k=2$ and $X^{\prime}$, there exists a graph $X^{\prime\prime}$ with $\operatorname{rtd}_{2}(X^{\prime\prime})=\operatorname{rtd}_{2}(X^{\prime})=t$ such that for every covering $S_{1},S_{2}$ of $V(X^{\prime\prime})$ there exists $i\in\{1,2\}$ such that $X^{\prime\prime}$ contains an $\{H\subseteq X^{\prime\prime}\mid\text{$H$ connected and }S_{i}\cap V(H)\neq\emptyset\}$-rich model of $X^{\prime}$. Let $\bar{c}_{1}(t,Y)=2+\bar{c}(t-1,X^{\prime\prime})$.
@@
Let $\mathcal{F}^{\prime}$ be the family of all the connected subgraphs $H$ of $G-U$ such that $U\cap N_{G}(V(H))\neq\emptyset$ and $F\subseteq H$ for some $F\in\mathcal{F}$. We argue that there is no $\mathcal{F}^{\prime}$-rich model of $X^{\prime\prime}$ in $G-U$. Suppose to the contrary that there is an $\mathcal{F}^{\prime}$-rich model $\big{(}B_{x}\mid x\in V(X^{\prime\prime})\big{)}$ of $X^{\prime\prime}$ in $G-U$. For each $u\in U$, let $S_{u}=\{x\in V(X^{\prime\prime})\mid u\in N_{G}(B_{x})\}$. Since the model is $\mathcal{F}^{\prime}$-rich, $\{S_{u}\}_{u\in U}$ is a covering of $V(X^{\prime\prime})$. Therefore, there exists $u\in U$ such that $X^{\prime\prime}$ contains an $\{H\subseteq X^{\prime\prime}\mid\text{$H$ connected and }S_{u}\cap V(H)\neq\emptyset\}$-rich model $\mathcal{M}$ of $X^{\prime}$. Moreover, by definition, each branch set of $\mathcal{M}$ contains a member of $\mathcal{F}$, and furthermore, $\mathcal{M}$ is a model of $X^{\prime}$ in $G$. Recall that $X^{\prime}=K_{1}\sqcup Y$. If $v$ is the vertex of $K_{1}$ is $X^{\prime}$, then the model obtained from $\mathcal{M}$ by replacing the branch set $C_{v}$ corresponding to $v$ by $C_{v}\cup\{u\}$ yields an $\mathcal{F}$-rich model of $K_{1}\oplus Y$ in $G$, which is a contradiction. This proves that there is no $\mathcal{F}^{\prime}$-rich model of $X^{\prime}$ in $G-\{u\}$.
@@
    for every $x\in V(T_{C})$,
    $\operatorname{wcol}_{r}\left(G_{C,x},P_{C,x}\right)\leqslant\bar{c}_{1}(t,Y)k\cdot r^{t-3}\log r$
    for every integer $r$ with $r\geqslant 2$, where, for $T_{C,x}$ being the subtree of $T_{C}$ rooted in $x$, $G_{C,x}$ is the subgraph of $G_{C}$ induced by $U_{C,x}=\bigcup_{y\in V(T_{C,x})}P_{C,y}$ and the vertex sets of all the components of $G_{C}-S_{C}$ having a neighbor in $U_{C,x}$.
@@
Since for every $C\in\mathcal{C}_{1}$, $U\cap N_{G}(V(C))=\emptyset$, every edge in $G[S]$ containing a vertex in $U$ has another endpoint in $U\cup S_{0}=P_{s}\cup P_{s^{\prime}}$. Consider an edge $vw$ in $G[S]$ such that $v\in S_{0}$ and $w\in S_{C}$ for some $C\in\mathcal{C}_{1}$. Since $\big{(}T_{C},(P_{C,x}\mid x\in V(T_{C}))\big{)}$ is a tree partition of $G_{C}[S_{C}]$ with $P_{C,s_{C}}=U_{C}$ and $S_{0}\subseteq V((G-U)-V(C))$, we conclude that $w\in P_{x}$ for some $x\in V(T_{C})$ such that $s^{\prime}x$ is an edge in $T$. Finally, for every edge $vw$ of $G[S]$ with $v,w\not\in U\cup S_{0}$, $vw$ is an edge of $G[S_{C}-U_{C}]$ for some component $C\in\mathcal{C}_{1}$, and so $v\in P_{C,x}$ and $w\in P_{C,y}$ for adjacent or identical vertices $x,y$ of $T_{C}$. Then $v\in P_{x}$ and $w\in P_{y}$. It follows that $\big{(}T,(P_{x}\mid x\in V(T))\big{)}$ is a tree partition of $G[S]$.
@@
Finally, we argue (d). Let $r$ be an integer with $r\geqslant 2$ and let $x\in V(T)$. For $x=s$, $|P_{s}|=|U|\leqslant 2$, thus the assertion is clear. For $x=s^{\prime}$, we have that $G_{s^{\prime}}$ is a union of components of $G-U$. By (C’),
$\operatorname{wcol}(G_{s^{\prime}},S_{0})=\operatorname{wcol}_{r}(G-U,S_{0})\leqslant\bar{c}(t-1,X^{\prime\prime})k\cdot r^{t-3}\log r\leqslant\bar{c}_{1}(t,Y)k\cdot r^{t-3}\log r.$
For $x\in V(T_{C}-\{s_{C}\})$ for some $C\in\mathcal{C}_{1}$, we have $G_{x}=G_{C,x}$, thus, the asserted inequality follows from (d”). This ends the proof of the claim. ∎
@@
Let $Y$ be a graph with $\operatorname{rtd}_{2}(Y)\leqslant t-1$. There are integers $\bar{c}_{2}(t,Y)$ such that for every positive integer $k$, for every graph $G$ with $\operatorname{tw}(G)<k$, for every family $\mathcal{F}$ of connected subgraphs of $G$, if $G$ has no $\mathcal{F}$-rich model of $K_{1}\oplus Y$, then there exist $S\subseteq V(G)$ such that
@@
Proof of the claim.
Let $\bar{c}_{2}(t,Y)=2\cdot\bar{c}_{1}(t,Y)$ and let $r$ be an integer with $r\geqslant 2$. We apply 4 with an arbitrary singleton $\{u\}$ in $G$ to obtain $S\subseteq V(G)$, a tree $T$ rooted in $s\in V(T)$, and a tree partition $(T,(P_{x}\mid x\in V(T)))$ of $S$ in $G$ with $P_{s}=\{u\}$ such that
@@
    for every $x\in V(T)$,
    $\operatorname{wcol}_{r}\left(G_{x},P_{x}\right)\leqslant\bar{c}_{1}(t,Y)k\cdot r^{t-3}\log r$
    for every integer $r$ with $r\geqslant 2$, for $T_{x}$ being the subtree of $T$ rooted in $x$, $G_{x}$ is the subgraph of $G$ induced by $\bigcup_{y\in V(T_{x})}P_{y}$ and the vertex sets of all the components of $G-S$ having a neighbor in $U_{x}$.
@@
$|\operatorname{WReach}_{r}[G,S,\sigma,u]\cap P_{y}|\leqslant\operatorname{wcol}_{r}\left(G_{y},P_{y}\right)\leqslant\bar{c}_{1}(t,X)k\cdot r^{t-2}\log r.$
@@
This concludes the proof of (c), and the claim follows. ∎
@@
Let $Y$ be a graph with $\operatorname{rtd}_{2}(Y)\leqslant t-1$ and let $h,d$ be positive integers. There exists an integer $\bar{c}_{3}(Y,h,d)$ such that for every positive integer $k$, for every graph $G$ with $\operatorname{tw}(G)<k$, for every family $\mathcal{F}$ of connected subgraphs of $G$, if $G$ has no $\mathcal{F}$-rich model of $T^{\prime}_{h,d}(Y)$, then there exists $S\subseteq V(G)$ such that
@@
Proof of the claim.
We proceed by induction on $h$. When $h=1$, $T^{\prime}_{h,d}(Y)=K_{1}\oplus(Y\sqcup Y)$ and the result follows from the previous claim applied to $Y\sqcup Y$ (note that $\operatorname{rtd}_{2}(Y\sqcup Y)=\operatorname{rtd}_{2}(Y)\leqslant t-1$) by setting $\bar{c}_{3}(Y,1,d)=\bar{c}_{2}(t,Y\sqcup Y)$.
@@
Finally, by Lemma 28, for every for every graph $X$ with $\operatorname{rtd}_{2}(X)\leqslant t$, there exists a graph $Y$ with $\operatorname{rtd}_{2}(Y)\leqslant t-1$ and positive integers $h,d$ such that $X\subseteq T_{h,d}(Y)\subseteq T^{\prime}_{h,d}(Y)$. Let $G$ be a graph with $\operatorname{tw}(G)<k$ and let $\mathcal{F}$ be a family of connected subgraphs of $G$. Suppose that $G$ has no $\mathcal{F}$-rich model of $X$. By 6, there exists $S_{0}\subseteq V(G)$ such that
@@
Item (A) follows from 6.(a). Let $C^{\prime}$ be a component of $G-S$, and let $C$ be the component of $G-S_{0}$ intersecting $V(C^{\prime})$. Then $N_{G}(V(C^{\prime}))\subseteq N_{G}(V(C))\cup\bigcup_{Q\in\mathcal{Q}_{C}}V(Q)$, and since $\bigcup_{Q\in\mathcal{Q}_{C}}V(Q)$ induces a connected subgraph of $C$ and contains a neighbor of every component of $G-V(C)$ having a neighbor in $V(C)$, we deduce that $N_{G}(V(C^{\prime}))$ intersects at most one component of $G-V(C^{\prime})$. This proves that (B) holds. And finally, (C) is true once we set $\bar{c}(t,X)=\bar{c}_{3}(Y,h,d)+(4h-1)\cdot 3$ since